OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7

OpenStudy (anonymous):

is this multiple choice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you type out the choices so i can graph them and see which one passes through (-2,2) ?

OpenStudy (anonymous):

@jocelynsalcedo :)

OpenStudy (anonymous):

y = 4x + 10 y = 1/4x + 5/2 y = -4x - 6 y = -1/4x + 3/2

OpenStudy (nurali):

y = 4x + 7 we know that the slope of a parallel line must be of the form: y = 4x + b Next the problem tells us that the parallel line we are looking for goes through the point (-2, 2). That means that when x = -2 and y = 2, the equation must be true. So we can take our form: . y = 4x + b . and substitute -2 for x and 2 for y and then find the value of b. Making these two substitutions for x and y, we transform the equation to: . 2 = 4*-2 + b . The 4*-2 multiplies out to -8 which makes the equation: 2 = -8 + b 10 = b y = 4x + 10

OpenStudy (anonymous):

thanks.

OpenStudy (nurali):

Anytime.